A couple purchasing a home budget $1600 per month for their loan payment. If they have $30,000 available for a down payment and are considering a 25-year loan, how much can they spend on the home at each of the following rates? (Round your answers to the nearest cent.)

(a)    6.3% compounded monthly
(b)    7.1% compounded monthly
Answer #1
Req a:
Annual rate =6.3%
Monthly rate = 6.3 / 12 = 0.525%
Monthly payment 1600
Multiply: Annuity PVF at 0.525% for 300 periods 150.8834
Present value of payment 241413.4
Add: Down payment 30000
Maximum amount that can be spend 271413.4
Req b:
Annual rate = 7.1%
Monthly rate = 7.1 /12 = 0.5917%
Monthly payment 1600
Multiply: Annuity PVF at 0.5917% for 300 periods 140.2137
Present value of payment 224342
Add: Down payment 30000
Maximum amount that can be spend 254342